OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x − 3y = 3 5x − 4y = 3 x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system −3 3

OpenStudy (anonymous):

4x-3y=3 /*4 16x-12y=12 A 5x-4y=3 /*3 15x-12y=9 B Subtract B from A to get x=3